ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 41, 60 Dȩbicki, Krzysztof, University of Wrocław, Poland Asymptotic of the tail distribution of supremum from a Gaussian process Let X(t) be centered Gaussian process with stationary increments. We use double sum method to obtain the exact symptotic of P(sup_{t≥0}(X(t)−t) > u) for u → ∞. In particular we consider X(t) = ∫_{0}^{t} Z(s)ds, where {Z(s) : s ≥ 0} is stationary centered Gaussian process. These results are joint work with prof. Tomasz Rolski. Reference: Dȩbicki, K., Rolski, T. (1998) An exact asymptotic of the tail distribution of supremum form a Gaussian process. Technical Report, Mathematical Institute, Wrocław University.